f(x)=(sinx+cosecx)2+(cosx+secx)2 then find minimum value of f(x)

Grade:12

1 Answers

Khimraj
7 years ago
Just open the expression.F(x) = sin²x + cosec²x + 2 + cos²x + sec²x +2,So F(x) = 5+ (sin²x +cos²x)/sin²xcos²x,Then F(x) = 5 + 4/sin²2x,Now you can easily see than max value of sin²2x is 1 and Min value of F(x) is 9.Hope it clears. If still u have then please clarify. If u like answer then please approve the answer.